# STEMS: Spatial-Temporal Enhanced Safe Multi-Agent Coordination for Building Energy Management

### v1

@@ III-A3 Evaluation Metrics
Discomfort Proportion: Percentage of time when indoor temperature is outside the comfort range [20°C, 26°C]:
$D_{prop}=\frac{1}{T\cdot N}\sum_{t=1}^{T}\sum_{i=1}^{N}\mathbf{1}(T_{i,t}<20\text{ or }T_{i,t}>26)$
where $T_{i,t}$ is the indoor temperature of building $i$ at time $t$.
Safety Violations(%): Percentage of safety constraint violations including battery SOC limits, thermal comfort bounds, and equipment operation constraints.
All metrics except Safety Violations are normalized relative to the rule-based baseline performance to enable fair comparison across different methods and scenarios.
@@ III-B Performance Evaluation
TABLE I: Long-term Performance Comparison (12-Month Average Results)
Peak Consumption Rate Proportion Violations(%)
Rule-Based 1.00 1.00 1.00 1.00 1.00 0.13 35.1
MPC 0.87 0.91 0.98 0.97 0.98 0.65 33.0
Single-Agent SAC 0.82 0.86 0.85 0.81 0.92 0.48 22.3
MADDPG 0.79 0.81 0.85 0.85 0.93 0.35 19.7
MetaEMS 0.83 0.80 0.88 0.83 0.90 0.39 23.1
MARLISA 0.84 0.87 0.82 0.83 0.95 0.23 21.4
STEMS 0.79 0.82 0.82 0.80 0.88 0.13 5.6
@@
[…]
Over the 12-month evaluation period, STEMS demonstrates sustained superior performance across almost all metrics compared to baseline methods. Traditional control approaches exhibit significant limitations in economic aspects in multi-building coordination. Rule-based control also suffers from rigid scheduling policies that cannot adapt to dynamic conditions, resulting in the highest safety violation rate (35.1%). MPC, while offering some optimization capability, is limited by its reliance on simplified linear models and lacks coordination mechanisms, achieving only modest improvements with 33.0% safety violations and high discomfort proportion. For learning-based baselines, Single-Agent SAC, despite showing competitive economic performance (0.82 cost, 0.86 emission), suffers from lack of coordination and safety guarantees, resulting in 22.3% safety violations and suboptimal comfort management (0.48 discomfort proportion). MADDPG and MetaEMS, while incorporating multi-agent learning, lack explicit safety mechanisms and spatial-temporal modeling, leading to higher safety violation rates (19.7% and 23.1% respectively) and limited comfort optimization. MARLISA achieves reasonable performance but still exhibits 21.4% safety violations due to absence of safety constraints. STEMS’s integrated approach combining spatial-temporal graph representation with CBF-based safety mechanisms delivers superior overall performance, while maintaining the lowest safety violation rate (5.6%) and best comfort management (0.13 discomfort proportion).
Then we evaluate the robustness and adaptability of STEMS under challenging conditions over the short-term control, we conduct detailed case studies during specific 2-3 day extreme weather. We select Extreme Weather Scenarios where Heat Wave: Outdoor temperature over 35-40°C (occurring 21.6% of the year), cooling demand increased by 150%, solar generation enhanced by 20%. Cold Wave: Outdoor temperature under -3°C (occurring 3.0% of the year), heating demand increased by 180%, battery capacity reduced by 15%.
Figure 3: Performance comparison between Heat Wave and Cold Wave scenarios. The charts display improvement rates relative to Rule-Based baseline across five key technical metrics: Cost Reduction, Emission Reduction, Peak Reduction, Consumption Reduction, and Ramping Reduction. STEMS (highlighted in red) demonstrates consistent superior performance in both extreme weather conditions, achieving 19% cost reduction and 16% peak reduction in Heat Wave scenarios, and 18% cost reduction with 16% peak reduction in Cold Wave scenarios.
Figure 4: Discomfort levels comparison between Heat Wave and Cold Wave scenarios across all evaluated methods. STEMS achieves the lowest discomfort levels in both scenarios (0.15 for Heat Wave, 0.19 for Cold Wave), representing 21% and 32% improvement over Rule-Based control respectively. The consistent low discomfort levels demonstrate STEMS’s superior thermal comfort management capabilities under extreme weather conditions.
Figure 5: Safety violations comparison between Heat Wave and Cold Wave scenarios. STEMS maintains the lowest safety violation rates (9.6% for Heat Wave, 8.6% for Cold Wave), representing 77% and 81% reduction compared to Rule-Based control. The CBF-based safety mechanism proves particularly effective in extreme conditions, ensuring operational safety while maintaining performance optimization.
Figure 3 presents a comprehensive performance analysis through radar chart visualization, revealing STEMS’s consistent superiority across both extreme weather scenarios. The radar charts display improvement rates relative to Rule-Based baseline, providing intuitive comparison of multi-dimensional performance metrics. In Heat Wave conditions, STEMS achieves remarkable technical improvements: 19% cost reduction, 16% emission reduction, 16% peak demand reduction, 18% consumption reduction, and 10% ramping reduction. The Cold Wave scenario demonstrates similar excellence with 18% cost reduction, 14% emission reduction, 16% peak reduction, 16% consumption reduction, and 11% ramping reduction. The consistent radar chart patterns across both scenarios highlight STEMS’s robust adaptability, with the spatial-temporal coordination mechanism effectively handling the contrasting challenges of cooling-dominated Heat Wave periods and heating-intensive Cold Wave conditions.
Heat Wave Case Study: During extreme heat conditions (over 35°C outdoor temperature), cooling demand increases by 150% while solar generation is enhanced by 20%, creating significant challenges for energy management systems. The radar chart analysis in Figure 3 demonstrates STEMS’s exceptional technical performance, achieving balanced improvements across all metrics. The spatial-temporal coordination enables intelligent scheduling of cooling systems across buildings, avoiding simultaneous activation that would overload the grid. The enhanced solar generation is effectively coordinated across buildings through the spatial graph network, enabling optimal energy sharing and storage scheduling during peak cooling periods. Most notably, as shown in Figure 4, thermal comfort is dramatically improved with STEMS achieving 0.15 discomfort level compared to 0.19 for Rule-Based control (21% improvement), while Figure 5 reveals safety violations are reduced by 77% (from 42.4% to 9.6%).
Cold Wave Case Study: Under severe cold conditions (below -3°C), heating demand surges by 180% while battery capacity decreases by 15% due to low-temperature effects. The radar chart visualization shows STEMS’s adaptive strategy, maintaining strong performance across all technical metrics despite the challenging operating environment. The algorithm automatically compensates for reduced battery performance by optimizing heating schedules and leveraging inter-building coordination through the learned spatial connections. Figure 4 shows thermal comfort improvement with STEMS achieving 0.19 discomfort level (32% improvement over Rule-Based), while Figure 5 demonstrates safety violations reduced by 81% (from 45.1% to 8.6%). The CBF-based safety mechanism proves particularly valuable in cold conditions, maintaining critical battery and thermal constraints despite the challenging operating environment.
The comprehensive visualization analysis demonstrates STEMS’s superior adaptability and robustness across diverse challenging conditions. The radar chart patterns reveal that STEMS maintains high performance levels even when individual components (batteries, HVAC systems) operate under stress, showcasing the value of multi-agent coordination and safety-constrained learning in real-world deployment scenarios. The consistent performance advantage across both extreme weather scenarios, combined with exceptional comfort management and safety assurance, validates STEMS’s practical viability for real-world building energy management systems operating under climate variability and extreme weather events.
@@ III-C System Visualization
To intuitively understand the spatial and temporal relationships learned by STEMS, in this subsection, we inspect the learned building attention weight matrix and temporal attention. We randomly select a set of buildings which includes 5 residential, 2 commercial and 1 mixed buildings . The building connection weight heatmap (Figure 6) reveals the learned spatial relationships among the 8 buildings in the multi-building energy management system. The visualization demonstrates a heterogeneous connection pattern with connection weights ranging from 0.008 to 0.731, reflecting STEMS’s ability to identify coordination opportunities and selective communication strategies. The strongest connections emerge between Buildings 3-4 (0.731) and Buildings 1-4 (0.614), indicating critical coordination pairs within the residential cluster that enable effective load balancing and energy sharing.
Residential buildings (Buildings 1-5) show varied intra-type connections with moderate-to-strong relationships, demonstrating selective coordination rather than uniform clustering. Commercial buildings (Buildings 6-7) exhibit moderate inter-connection (0.387), while the mixed-use building (Building 8) serves as a strategic coordination hub with balanced connections to both building types. Cross-type connections reveal strategic coordination patterns, with notable residential-commercial links such as Buildings 1-7 (0.324) and 4-7 (0.312) enabling load balancing during complementary demand periods. The learned connection weights automatically adapt to changing conditions, strengthening connections between buildings that benefit most from coordination while reducing unnecessary communication overhead.
@@
Figure 7: Building-type-specific temporal attention patterns comparison: (a) Residential building showing distributed attention with evening peaks and lifestyle-driven variability, (b) Commercial building exhibiting concentrated work-hour attention with high regularity and minimal after-hours activity.
The residential building temporal attention pattern (Figure 7a) exhibits distinct lifestyle-driven characteristics reflecting typical household energy consumption behaviors. The visualization shows attention weights ranging from 0.05 to 0.15, with distributed attention during morning hours and peak self-attention at 8:00 PM (0.18-0.20), corresponding to residential energy consumption peaks when families engage in evening activities.
In contrast, the commercial building attention pattern (Figure 7b) demonstrates highly structured temporal dependencies aligned with business operations. The 12:00 PM period exhibits the strongest attention weights (0.15-0.20) during peak business hours, while 8:00 PM shows dramatically reduced attention weights (0.02-0.05), reflecting minimal energy activity after business hours when buildings operate in energy-saving mode.
The complementary nature of these temporal patterns—with residential buildings showing peak activity during evening hours when commercial buildings are inactive—creates optimal opportunities for energy sharing and load balancing through the spatial connections. This spatial-temporal integration demonstrates STEMS’s ability to learn and leverage both the physical similarities between buildings and their temporal operational patterns, enabling load shifting and peak shaving through the spatial network while respecting the natural characteristics of different building types.
III-D Computational Complexity and Scalability
We then report the computational efficiency and scalability of STEMS across different building network sizes. Table II presents the relationship between building count, model size, and convergence time for scales ranging from 5 to 15 buildings. These scales represent realistic deployment scenarios commonly found in smart building systems, from small residential complexes to medium-sized neighborhoods.
TABLE II: Model Complexity and Scalability Analysis
Building Count Model Size Convergence Time
5 buildings 1.51 MB 28 min
8 buildings 1.91 MB 45 min
15 buildings 3.22 MB 158 min
@@ III-D Computational Complexity and Scalability
The model size scales approximately linearly with building count, primarily due to the independent actor-critic networks required for each building agent. The shared spatial-temporal graph components maintain relatively constant size, contributing to computational efficiency. The model size increases from 1.51 MB for 5 buildings to 1.91 MB for 8 buildings, demonstrating efficient scaling with manageable memory requirements. Convergence time exhibits quadratic growth $O(N^{2})$, reflecting the computational complexity of spatial-temporal graph learning where message passing scales with the square of building count. Despite this quadratic scaling, absolute convergence times remain practical for real-world deployment. The convergence time increases from 28 minutes for 5 buildings to 158 minutes for 15 buildings, which is acceptable for offline training.
III-E Ablation Study
Table III analyzes the contribution of each component in the STEMS framework.
TABLE III: Ablation Study Results
@@ III-E Ablation Study
Peak Consumption Rate Proportion Violations(%)
w/o GCN-Transformer 0.83 0.87 0.86 0.82 0.93 0.28 12.4
w/o Spatial Graph 0.81 0.85 0.84 0.81 0.91 0.23 9.8
w/o Temporal Attention 0.80 0.84 0.83 0.81 0.90 0.19 8.7
w/o CBF Safety 0.78 0.80 0.82 0.79 0.89 0.14 20.8
STEMS 0.79 0.82 0.82 0.80 0.88 0.13 5.6
@@
[…]
The most critical finding is CBF’s role as the primary safety differentiator. Removing CBF increases safety violations from 5.6% to 20.8%—crossing the threshold from acceptable operational risk to unacceptable system reliability. In real building management, this 271% increase represents the difference between regulatory compliance and potential equipment damage or occupant safety concerns. Notably, CBF achieves this safety improvement with minimal performance cost, maintaining nearly identical economic metrics while dramatically reducing operational risk. The GCN-Transformer architecture serves as the performance foundation, with its removal causing the largest individual degradation (5.1% cost increase, 115% comfort degradation). This reflects the practical reality that without proper spatial-temporal feature representation, coordination mechanisms cannot function effectively. The architecture’s impact on safety violations (121% increase to 12.4%) demonstrates that effective feature learning is prerequisite for both performance optimization and constraint satisfaction in complex multi-building systems. The spatial graph and temporal attention components demonstrate diminishing returns typical of real-world coordination systems. Spatial coordination provides moderate benefits (2.5% cost reduction, 77% comfort improvement from 0.23 to 0.13 discomfort proportion), reflecting the practical challenges of inter-building energy sharing due to transmission losses and grid constraints. Temporal attention shows smaller individual impact (1.3% cost reduction) but proves essential for predictive control, enabling proactive rather than reactive energy management—a crucial capability during peak demand periods.

### v2

@@ III-A3 Evaluation Metrics
Discomfort Rate: Proportion of occupied timesteps when indoor temperature deviates from setpoint by more than 2°C:
$D_{rate}=\frac{1}{T_{occ}}\sum_{t=1}^{T}\sum_{i=1}^{N}\mathbf{1}(O_{i,t}>0)\cdot\mathbf{1}(|T_{i,t}-T_{i,t}^{set}|>2)$
where $T_{occ}$ is the total number of occupied timesteps across all buildings, $T_{i,t}^{set}$ is the setpoint temperature of building $i$ at time $t$, and $O_{i,t}$ indicates occupancy status.
Safety Viol. Rate: Proportion of timesteps violating safety constraints.
All metrics except Discomfort Rate and Safety Violations Rate are normalized relative to the rule-based baseline performance to enable fair comparison across different methods and scenarios.
Figure 3: Performance comparison between Heat Wave and Cold Wave scenarios across five key metrics: Cost Reduction, Emission Reduction, Peak Reduction, Consumption Reduction, and Ramping Reduction. STEMS (red) achieves 18-19% cost reduction and 16% peak reduction in both scenarios.
@@ III-B Performance Evaluation
We also conduct case studies during extreme weather periods. Heat Wave conditions where temperature $\geq$ 30°C, occurring 14.8% of the year, create high cooling demand with enhanced solar generation. Cold Wave conditions where temperature $\leq$ 0°C, occurring 1.2% of the year, require substantial heating with minimal solar availability. These scenarios test system performance when individual components (batteries, HVAC systems) operate under stress.
Figure 3 presents performance analysis under extreme weather conditions. The radar charts display improvement rates relative to Rule-Based baseline across five key metrics, providing intuitive comparison across multiple dimensions. Under Heat Wave conditions, STEMS achieves 19% cost reduction, 16% emission reduction, 16% peak reduction, 18% consumption reduction, and 10% ramping reduction. The Cold Wave scenario demonstrates similar performance with 18% cost reduction, 14% emission reduction, 16% peak reduction, 16% consumption reduction, and 11% ramping reduction. The consistent radar patterns across both scenarios validate robust performance under contrasting thermal challenges. Figure 4 reveals distinct comfort management strategies. STEMS maintains the lowest discomfort rates with 0.153 for Heat Wave and 0.192 for Cold Wave, closely matched by D-MAPPO with 0.184 and 0.202. The tight comfort control demonstrates effective thermal management even when HVAC systems operate near capacity limits during extreme temperatures. Figure 5 shows safety performance differences. STEMS achieves the lowest violation rates with 0.096 for Heat Wave and 0.086 for Cold Wave, representing 77-81% reduction compared to Rule-Based control. MADCQ achieves moderate safety with 0.225 and 0.218 through heuristic constraint handling, while D-MAPPO shows higher violations with 0.283 and 0.275 despite good comfort performance. The performance gap highlights that penalty-based approaches and heuristic methods cannot provide the formal safety guarantees achieved by CBF mathematical optimization, especially critical when systems operate under stress during extreme weather events. STEMS’s spatial-temporal coordination enables effective energy management by optimizing battery reserves and coordinating inter-building information sharing before extreme events or during peak stress. Additionally, CBF safety guarantees maintain constraint satisfaction even when systems operate at their limits during simultaneous peak demand. In contrast, other baselines’ constraint handling approaches fail to provide formal guarantees under stress.
Figure 4: Discomfort rate comparison between Heat Wave and Cold Wave scenarios. STEMS maintains the lowest discomfort rate across all evaluated methods.
Figure 5: Safety violation rates comparison between Heat Wave and Cold Wave scenarios. STEMS achieves the lowest safety violation rates across all evaluated methods.
@@ III-C System Visualization
To intuitively understand the spatial and temporal relationships learned by STEMS, in this subsection, we inspect the learned building attention weight matrix and temporal attention. We randomly select a set of buildings which includes 5 residential, 2 commercial and 1 mixed buildings . The building connection weight heatmap in Figure 6 reveals the learned spatial relationships among the 8 buildings in the system. The visualization demonstrates a heterogeneous connection pattern with connection weights ranging from 0.008 to 0.731, reflecting STEMS’s ability to identify coordination opportunities and selective communication strategies. The strongest connections emerge between Buildings 3-4 with 0.731 and Buildings 1-4 with 0.614, indicating critical coordination pairs within the residential cluster that enable effective load balancing through information sharing. Residential buildings show varied connections with moderate to strong relationships, demonstrating selective coordination rather than uniform clustering. Commercial buildings exhibit moderate connection at 0.387, while the mixed building serves as a strategic coordination hub with balanced connections to both building types. Cross-type connections reveal strategic coordination patterns, with notable links such as Buildings 1-7 at 0.324 and 4-7 at 0.312 enabling load balancing during complementary demand periods. The learned connection weights automatically adapt to changing conditions, strengthening connections between buildings that benefit most from coordination while reducing unnecessary communication overhead.
@@
Figure 7: Building-type-specific temporal attention patterns comparison: (a) Residential building showing distributed attention with evening and morning peaks, (b) Commercial building exhibiting concentrated work-hour attention with high regularity and minimal after-hours activity.
The residential building temporal attention pattern in Figure 7a exhibits distinct characteristics reflecting typical household energy consumption behaviors. The visualization shows distributed attention during morning hours and peak attention at 8:00 PM, corresponding to residential energy consumption peaks when families engage in evening activities. In contrast, the commercial building attention pattern in Figure 7b demonstrates highly structured temporal dependencies aligned with business operations, reflecting minimal energy activity after business hours when buildings operate in energy saving mode.
The complementary nature of these temporal patterns creates opportunities for coordinated load balancing through information sharing via the spatial connections. Residential buildings show peak activity during evening hours when commercial buildings are inactive. This integration demonstrates STEMS’s ability to learn and leverage both the physical similarities between buildings and their temporal operational patterns, enabling load shifting and peak shaving through the spatial network while respecting the natural characteristics of different building types.
III-D Robustness Analysis
In this subsection, we evaluate the robustness of STEMS through: 1) power outage scenarios to test resilience under unexpected grid failures, and 2) the cold-humid Vermont dataset consisting of same-type residential buildings to evaluate cross-climate generalizability.
III-D1 Power Outage Scenarios
Power outages represent critical stress tests for building energy management systems, particularly during extreme weather when grid reliability decreases. We simulate stochastic power outage scenarios in CityLearn environment with configurations based on industry-standard reliability indices from U.S. EIA data: SAIFI=1.436 events/year and CAIDI=331.2 minutes/event (approximately 5.5-hour outages) [citylearn]. We evaluate system robustness under three weather conditions during which power outages occur: (1) Year-round: outages occur randomly throughout the year, (2) Heat Wave (T $\geq$ 30°C), and (3) Cold Wave (T $\leq$ 0°C). All methods were tested directly under outage conditions. Results averaged over 5 random seeds are presented in Table II. The unserved energy rate represents normalized energy demand unmet during outages (lower is better).
TABLE II: Power Outage Robustness under Different Weather Conditions.
Scenario Metric Rule- Single- MAD Meta MAD D- STEMS
Based Agent SAC DPG EMS CQ MAPPO
Year-round Unserved Energy Rate 0.784 0.663 0.605 0.578 0.582 0.554 0.553
Safety Viol. Rate 0.412 0.283 0.254 0.291 0.214 0.253 0.082
Heat Wave Unserved Energy Rate 0.813 0.705 0.642 0.618 0.624 0.601 0.594
Safety Viol. Rate 0.453 0.325 0.294 0.365 0.274 0.325 0.113
Cold Wave Unserved Energy Rate 0.842 0.736 0.674 0.648 0.653 0.635 0.623
Safety Viol. Rate 0.485 0.354 0.326 0.332 0.243 0.291 0.102
@@ III-D1 Power Outage Scenarios
Table II demonstrates STEMS’s strong robustness across diverse weather conditions during power outages. In the year-round baseline scenario, STEMS achieves 0.553 unserved energy rate, outperforming almost all comparison methods through GCN-based spatial coordination that enables load reduction strategies across buildings during outages. Most significantly, STEMS reduces safety violations to 0.082, representing 62-80% improvement over other methods, achieved through CBF-based mathematical safety guarantees that provide hard constraints under stress even without explicit power outage training. Under extreme weather conditions, the performance gap becomes more pronounced. During Heat Wave scenarios with high cooling loads and abundant solar generation, STEMS maintains 0.594 unserved energy rate and 0.113 safety violations. The Transformer’s temporal planning recognizes outage risk during extreme weather and maintains higher battery reserves through conservative charging strategies, indirectly preparing for potential outages. In the most challenging Cold Wave scenarios with minimal solar generation and high heating demands, STEMS achieves lowest 0.623 unserved energy rate and 0.102 safety violations. The GCN-based information sharing enables buildings to coordinate thermal management decisions based on neighbors’ battery status and thermal conditions, optimizing cluster-wide energy utilization during prolonged outages.
Despite CBF protection, STEMS maintains 0.082-0.113 residual safety violations during power outages. These occur during prolonged Cold Wave outages when battery capacity becomes insufficient to simultaneously maintain thermal comfort. In such cases, the CBF safety layer prioritizes hardware protection over temporary thermal discomfort, preventing catastrophic system failure.
III-D2 Cross-Climate Generalization
To evaluate STEMS’s generalization across different climate zones, we conduct experiments on the Vermont, Chittenden County dataset from CityLearn benchmark. This dataset exhibits cold-dominant characteristics with 23.6% Cold Wave days (T $\leq$ 0°C) and only 1.8% Heat Wave days (T $\geq$ 30°C). The temperature ranges from -28.3°C to 35.6°C with mean 9.1°C, representing a significantly different climate profile from the Texas dataset. The dataset contains 47 residential buildings with one year of hourly data (8,760 hours). We randomly selected 8 buildings and evaluated 12-month average results. Figure 8 presents STEMS’s performance compared with four representative learning-based methods: Single-Agent SAC, MADDPG, and the two newly added baselines MADCQ and D-MAPPO. MetaEMS and MARLISA show similar performance patterns to MADDPG and are omitted for clarity.
Figure 8: Vermont dataset performance comparison. STEMS achieves the best overall performance in cold-dominant climate.
From Figure 8, STEMS achieves 0.812 cost, 0.843 peak, and 0.821 consumption, outperforming the second-best MADDPG by 2.4%. For operational quality, STEMS achieves 0.862 ramping rate and 0.122 discomfort rate. Most importantly, STEMS reduces safety violations to 0.048, representing 70% improvement over MADCQ and 77% over D-MAPPO, validating that CBF-based guarantees remain effective across different climate zones.
Comparing Vermont with Texas reveals adaptation patterns. Energy efficiency shows 2-3% higher cost in Vermont due to reduced load diversity. The Vermont dataset consists entirely of residential buildings with similar occupancy patterns, resulting in synchronized peak demands that limit peak shaving opportunities through temporal coordination. In contrast, Texas’s mixed commercial-residential portfolio provides complementary load profiles where GCN-based information sharing enables more effective coordination. However, ramping rate improves to 0.862 as the Transformer better learns consistent residential consumption patterns. Safety violations decrease to 0.048 since homogeneous buildings share uniform constraints, simplifying CBF optimization. Discomfort rate improves to 0.122 through more uniform comfort-energy trade-offs. These results validate STEMS’s cross-climate generalization capability and its adaptability to both heterogeneous and homogeneous building settings.
III-E Computational Complexity and Scalability
We then report the computational efficiency and scalability of STEMS, where Table III compares STEMS with MPC and the state-of-the-art D-MAPPO and presents STEMS’s results across 5 to 15 buildings. For 8 buildings, STEMS requires 45 minutes of offline training, comparable to D-MAPPO’s 41 minutes, while enabling months of continuous operation. STEMS’s online inference completes in 24 ms, well within the typical 1-hour control cycle for real-time building energy management. In contrast, MPC requires 2.3 seconds per decision due to QP solving, introducing potential latency in time-critical scenarios. Furthermore, we find that STEMS achieves better safety at matched computational cost. At identical 8-building scale with similar training time, STEMS achieves 0.056 safety violation rate versus D-MAPPO’s 0.198, representing a 72% reduction. This improvement stems from CBF-QP optimization that enforces hard constraints at each control step, adding less than 1 ms to the inference time. The result demonstrates that formal safety guarantees need not compromise computational efficiency.
We also evaluated scalability of STEMS across 5 to 15 buildings, representing realistic deployment scenarios such as university campuses, corporate complexes, and residential communities [tariq2025dmappo, safeMBRL, chen2024green]. Model size scales linearly with building count, from 1.51 MB to 3.22 MB, as shared spatial-temporal graph components remain constant while independent actor-critic networks grow proportionally. Training time follows $O(N^{2})$ complexity due to graph message passing, increasing from 28 to 158 minutes. Despite this quadratic scaling, convergence times remain acceptable for offline training, while inference maintains $O(1)$ complexity per building through fixed network forward passes.
TABLE III: Computational Efficiency and Safety Performance Comparison.
Method Building Training Inference Model Safety
Count Time Time Size Viol. Rate
MPC 8 N/A 2.3 s N/A 0.351
D-MAPPO 8 41 min 22 ms 1.18 MB 0.198
STEMS 5 28 min 17 ms 1.51 MB 0.055
8 45 min 24 ms 1.91 MB 0.056
15 158 min 30 ms 3.22 MB 0.061
[…]
III-F Ablation Study
Table IV analyzes the contribution of each component in the STEMS framework.
TABLE IV: Ablation Study Results
@@ III-F Ablation Study
Peak Consumption Rate Rate Viol. Rate
w/o GCN-Transformer 0.832 0.874 0.863 0.825 0.931 0.283 0.124
w/o Spatial Graph 0.813 0.856 0.842 0.817 0.912 0.234 0.098
w/o Temporal Attention 0.803 0.842 0.831 0.814 0.904 0.193 0.087
w/o CBF Safety 0.782 0.801 0.823 0.796 0.894 0.145 0.208
STEMS 0.792 0.824 0.821 0.805 0.883 0.132 0.056
@@
The most critical finding is CBF’s role as the primary safety differentiator. Removing CBF increases safety violations from 0.056 to 0.208, crossing the threshold from acceptable operational risk to unacceptable system reliability. In real building management, this 271% increase represents the difference between regulatory compliance and potential equipment damage or occupant safety concerns. Notably, CBF achieves this safety improvement with minimal performance cost, maintaining nearly identical economic metrics while dramatically reducing operational risk. The GCN-Transformer architecture serves as the performance foundation, with its removal causing the largest individual degradation with 5.1% cost increase and 115% comfort degradation. This reflects the practical reality that without proper feature representation, coordination mechanisms cannot function effectively. The architecture’s impact on safety violations shows 121% increase to 0.124, demonstrating that effective feature learning is prerequisite for both performance optimization and constraint satisfaction in complex systems. The spatial graph and temporal attention components demonstrate diminishing returns typical of real-world coordination systems. Spatial coordination provides moderate benefits with 2.5% cost reduction and 77% comfort improvement from 0.234 to 0.132 discomfort rate, reflecting the practical challenges of inter-building coordination due to transmission losses and grid constraints. Temporal attention shows smaller individual impact with 1.3% cost reduction but proves essential for predictive control, enabling proactive rather than reactive energy management during peak demand periods.